OpenStudy (anonymous):

Solve for x. Graph the solution 9x+2<6x-13

OpenStudy (tkhunny):

Have you considered 1) Subtracting 2 2) Subtracting 6x 3) Dividing by 3 ??

OpenStudy (anonymous):

No haven't tried that my main focus is trying to graph it but without know what is the y it is hard to use a graphing calculator online

OpenStudy (tkhunny):

Ah, well, do those three things to both sides if the inequality and see what remains.

OpenStudy (triciaal):

one approach graph the 2 lines first y =9 x + 2 and y = 6 x -13

OpenStudy (anonymous):

did that but not understanding the the graphs

OpenStudy (triciaal):

for each value of x on the line 6 x -13 compare with the value of y on the line 9 x + 2 when is the value of (9 x + 2) < 6 x -13

OpenStudy (triciaal):

from the calculation we already know 9 x + 2 = 6 x - 13 when x = -5

OpenStudy (triciaal):

example when x = 2 6*2-13 = -1 the y value at x = 2 on the line 9 x + 2 = 9*2 + 2 = 20 so 9 x + 2 is not less than 6 x - 13 when x = -8 6*-8 -13 = -48-13 = -61 the y value at x = -8 on the line 9 x + 2 = 9*-8 + 2 = -72 +2 = -70 and -70 < -61 so x = -8 is a solution do you understand the lines any better now?

OpenStudy (triciaal):

 this is a rough sketch for any value of x>-5 the y value on the graph 9 x+ 2 is more than 6 x -13